What is the slope and y-intercept of the points (-4,5) and (6,25)?
Answer by Cromlix(4381)

using gradient formula to find 'm'
y2-y1/x2-x1
25 - 5/6 -(-4)
= 20/10
m = 2
using formula y - b = m(x - a) and Coords (6,25)
y - 25 = 2(x - 6)
y - 25 = 2x - 12
y = 2x - 12 + 25
y = 2x + 13
Hope this helps.
:-)